OpenStudy (anonymous):

Are column transformations valid to do while finding solution to Ax = 0. Doesn't it modify the nullspace? In what situations is it okay to do column transformations?

OpenStudy (kutulu):

Column transformations are never valid when trying to find the null space. If you think about the matrix as a series of linear equations, a column transformation would change the coefficients of one variable in each equation, without changing the solution to that equation. The result is not the same equation you started with.

OpenStudy (anonymous):

When you perform column operations, you have to account for any transformation you are performing, because, as kutulu remarked, you are generally altering the solution: Ex: x+y=0 => x=-y transform second column e.g. scale by 2: x+2y=0 => y'=2y now: x=-2y' => x=-2(1/2)*y=-y Interchange columns one and two: x'+y'=0, x'=y y'=x Similarly, you'd have to account for replacing column k by (column k) + (a*column j)